w minus negative 3 equals negaative 2, what is w?
Answer by jim_thompson5910(35256)

w - (-3) = -2

w +3 = -2

w +3 - 3 = -2 - 3

w = -5


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim